# SPYGLASS. VII-A. The Demographics and Ages of Small Nearby Young Associations

### v1

@@ 4.1.3 Lithium
The depletion of the Li-6707Å line can be used not just for youth indication as discussed in Section 3.1.5, but also as a powerful tool for age measurement. The EAGLES package provides empirical models of the lithium depletion curve as a function of age, as well functions for computing an age probability posterior for a given stellar Li EW and effective temperature $T_{eff}$. These functions use both the original analytical model (v1; Jeffries 2023) that we used in Section 3.1.5, and a new machine learning (ML) neural network version (v2; Weaver et al. 2024).
@@
We multiply the age likelihoods of all remaining stars to produce aggregate posteriors for the entire population, and we do this separately for both the original and ML EAGLES modules. We set the age to the 50th percentile of the posterior, and set the lower and upper uncertainty intervals to the 16th and 84th percentiles, respectively. We exclude ages under 5 Myr, where EAGLES lacks age discrimination and therefore sets the likelihood to its value at 5 Myr (Jeffries, 2023). We only report an upper age limit for populations where the posterior at 5 Myr exceeds half of its peak, which we set to the 95th percentile of the probability posterior. We plot the Lithium depletion age solutions in Figure 8, showing the Li EWs used in the calculation, the best-fit EAGLES Li curves, and the resulting combined age posterior.
The two different EAGLES modules produce broadly similar results. The main difference is in the “Li dip”, which reaches zero about 10 Myr earlier and transitions more sharply analytical version. In populations like AqE, where this dip is well-covered, the analytical model is in close agreement with both the nearly Li-free stars in the dip and the Li-rich stars to the right of it, while the ML version does not capture this sharp transition as closely. However, the ML version seems to produce more consistent results in populations with less coverage, especially populations with a significant Li scatter. The large scatter and Li dip shape in the ML module are produced by an observed scatter in real Li observations (Weaver et al., 2024). Our harsh quality cuts on binarity and membership may improve this scatter beyond the expectations of the ML model in populations like AqE, but this does not appear to be universal. We therefore include both the analytical and ML results in Figure 8 and Section 4.1.4, but use only the analytical results for membership assessment and quality checks for consistency with previous work (e.g., SPYGLASS-V).
@@ 4.1.4 Final Ages
We have produced isochronal ages using several different models, as well as dynamical and lithium depletion ages. The isochronal ages provide a consistent relative age scale derived from the height of the PMS that is available for all populations, unlike the dynamical and lithium depletion methods. However, isochronal ages are often inconsistent with each other and with other age calculation methods in absolute terms (e.g., Herczeg & Hillenbrand, 2015; Rottensteiner & Meingast, 2024). The accuracy of these isochronal solutions can be assessed by comparison with the non-isochronal methods.
@@
While agreement with other methods may marginally favor the PARSEC over DSEP-Magnetic, the use of PARSEC is perhaps better motivated by the consistent agreement between that model and our pre-main sequences. In Figure 6, more luminous stars tend to fall below the best-fit DSEP-Magnetic isochrone We therefore conclude that PARSEC produces the most reliable age scale for our analysis.
@@
The ML module, which is designed to better capture natural scatter in these measurements, produces results that align more closely with PARSEC in the populations with high analytical Li ages relative to PARSEC. The apparent high bias in the Li age fits may therefore be driven by coverage that is too incomplete to reject unrealistically old age solutions.
All dynamical ages are consistent with our PARSEC ages within uncertainties, except in SCYA-79, where the PARSEC ages are poorly constrained. There is no clear evidence of an offset between the dynamical and isochronal ages like what has previously been used to indicate a gas dispersal timescale (e.g., Kerr et al., 2022b; Miret-Roig et al., 2024). We discuss this point further in Section 5.
@@ 4.2 Demographics
All of the populations we explore in this paper have yet to have their sizes, masses, and total membership comprehensively investigated. SPYGLASS-IV’s candidate member lists provided the first look at sizes of many of these associations, while some larger populations have also been covered by the (Kounkel & Covey, 2019) membership lists. However, these publications did not address all potential sources of bias. Most of these populations represent an understudied demographic of low-mass, isolated populations that has yet to be explored, so detailed demographic work is essential to identify differences and similarities between these populations and the much larger populations that dominate the current literature on young associations.
@@ 4.2.1 Stellar Masses
We compute the masses of individual stars by comparing the stellar photometry to the PARSEC v1.2S isomass track grid from SPYGLASS-V, which uses mass sampling every 0.005 $M_{\odot}$ for $0.09M_{\odot}<M<1M_{\odot}$, every 0.01 $M_{\odot}$ between $1.0M_{\odot}<M<2.0M_{\odot}$, every 0.02 $M_{\odot}$ for $2M_{\odot}<M<4M_{\odot}$, and every 0.05 $M_{\odot}$ for $4M_{\odot}<M<20M_{\odot}$. We set the mass to that of the nearest isomass track. Most populations have a maximum stellar mass between 2 and 4 $M_{\odot}$, with no candidates more massive than 4 $M_{\odot}$. These populations therefore lack stars that could constrain the main sequence turn-off ages, which is not surprising given their small total masses (e.g., Chabrier, 2005).
@@ 4.2.2 Correction for Low-Mass Stars
We show a histogram of stars that pass the astrometric and photometric quality flags in Fig. 10, with bins evenly log-sampled between 0.09 and 15 M, and with each object weighed by $P_{fin}$. We smooth the result with a savitsky-golay filter to produce a mass frequency curve, which we also show in Fig. 10, alongside the Chabrier (2005) IMF scaled to best fit the mass frequency curve for $M>0.25M_{\odot}$ via least-squares optimization. Our mass distributions generally follow the Chabrier (2005) IMF, although there is some variation, most notably a stellar deficit at $M\sim 0.5$ and an excess at $M\sim 0.25$. This pattern has been shown in several prior publications, and is usually attributed to model innacuracies that displace some stars to lower masses (e.g., Kraus et al., 2014; Kerr et al., 2024).
@@
Figure 10: IMF-based mass corrections to all 15 associations where demographics are possible. The dark blue histogram shows the distribution of association members, and a smoothed version of that is shown in light blue. The orange curve shows the Chabrier (2005) IMF. The missing mass, shown in red, is computed using the volume between the Chabrier (2005) IMF and the smoothed version of the stellar mass distribution curve. This is used to compute the corrective factors for mass and number ($C_{M,lm}$ and $C_{N,lm}$, respectively), which are annotated on each panel.
All mass frequency curves show fewer stars in the low-mass regime compared to the Chabrier (2005) IMF, indicating a drop in Gaia completeness there. We take the difference between the mass distribution curve and the scaled IMF to the left of the leftmost crossover point in Figure 10 as the volume missed by Gaia. The corrective factor to number of members is the sum across the entire stellar histogram, plus the sum over the area to the left of the crossover point, divided by the sum over the entire stellar histogram. We limit this calculation to stars with masses $M>0.09M_{\odot}$, or above the PARSEC lower mass limit, which is used as the mass cutoff for inclusion in our stellar sample. The mass corrective factor is computed the same way, but with each histogram and scaled IMF multiplied by the mass at each bin element. We did not bound this calculation in mass, as low-mass and sub-stellar objects can still contribute to the virial state of a stellar population. The resulting corrective factors range from 1.06 to 1.18 in mass and from 1.14 to 1.53 in number. The size of these corrections correlate with distance, with four of five populations with a mass correction less than 1.1 having d<220d<220pc. SCYA-54, which contains the most distant subgroup we cover here, has the largest correction in both mass and number.
@@ 4.2.3 High-mass stars
Some of our mass histograms in Fig. 10 show a deficit of stars with $M\gtrsim 2M_{\odot}$, which may be caused by a lack of quality Gaia astrometry and photometry for stars that saturate ($G\lesssim 3$) (Gaia Collaboration et al., 2016b). However, Hipparcos provides complementary coverage that is essentially complete for stars with $G\lesssim 7$ (ESA, 1997; van Leeuwen, 2007). We therefore search each population for Hipparcos stars missed by Gaia. We mark any Hipparcos star as a potential member if its distance to the tenth nearest association member in SPYGLASS-IV’s space-transverse velocity coordinate space ($d_{10}$) is less than that of the member with the greatest $d_{10}$, following previous SPYGLASS publications. We only consider stars without Gaia astrometry to exclude stars that have already had their membership assessed.
Across all 16 populations, we identify only 5 credible additional members in the Hipparcos catalog across 4 of our associations. Only 3 stars: Mirzam ($\beta$ CMa) and HIP 32848 in CMaN and HIP 4285 in SCYA-54 have masses in the $M\gtrsim 2M_{\odot}$ range where we see a stellar deficit in Fig. 10. We therefore conclude that there is no evidence of our Gaia-based sample systematically excluding high-mass stars, and choose not to apply a broad mass correction. It is possible that high-mass stars are intrinsically more difficult to detect, perhaps due to an increased velocity dispersion produced by the higher binarity rates of high-mass stars (Sullivan & Kraus, 2021). Follow-up work will therefore be necessary to determine whether some populations truly have mass distributions that differ from a standard IMF.
@@ 4.2.4 Binaries
We estimate the mass contribution from unresolved companions by applying literature binary statistics to our detected primaries. We first identify resolved binaries to avoid double-counting companions when applying binary statistics. Following SPYGLASS-V, we search a radius of $10^{4}$ AU on-sky around each star in Gaia. We mark stars with $\Delta v_{T}<3$ km s-1 and $\frac{\Delta\pi}{\pi}<0.2$ as in the same stellar system, with the brightest member marked as the primary, and dimmer members considered companions. These conditions roughly encapsulate the full range of expected velocities for stars near Gaia’s roughly 1 arcsecond separation limit, given typical uncertainties (Rizzuto et al., 2018). Unresolved companions move the photometry of a star largely perpendicular to the pre-main sequence, changing its mass estimate relatively little. For measuring demographics, probable unresolved binaries with RUWE>1.2>1.2 can therefore be treated as if they are system primaries without substantially increasing the systematic uncertainties in our association mass estimates.
We follow the methods applied in SPYGLASS-VI to estimate the mass in binaries, sampling $10^{4}$ stars from the Chabrier (2005) IMF at random for $0.09<M<4~M_{\odot}$, and computing the expected mass and number of companions for each sampled star according to the mean mass ratios from Sullivan & Kraus (2021) and the companion fractions from Sullivan & Kraus (2021) and Duchêne & Kraus (2013). The expected binary contribution computed from this sampling increases population mass by a factor of 1.23, and increases number of stars by a factor of 1.39. These values produce an average correction to the contribution from binaries for any given stellar primary, including low-mass stars that we correct for in Section 4.2.2.
@@
ASSOCaaUnique choices and properties for specific associations: in CasE, we set stellar membership probabilities below a cut in Fig. 6 to zero to exclude an elevated field sequence. TOR1A and TOR1B are components of the TOR1 association that have velocities inconsistent with common formation. We provide basic position and velocity data for SCYA-79, but do not report population size, mass, or any values deriving from them due to weak separation with the field. SC NbbFor our mass and stellar population measurements, we adopt a 10% systematic uncertainty. MbbFor our mass and stellar population measurements, we adopt a 10% systematic uncertainty. RA Dec l b d $\overline{\mu_{RA}}$ $\overline{\mu_{Dec}}$ $\overline{v_{T,l}}$ $\overline{v_{T,b}}$ $\overline{v_{r}}$ $R_{hm}$ $\sigma_{1D}$ $\sigma_{vir}$ Vir. Rat.
@@
AndS 69 30.2 6.4 30.3 116.3 -32.1 169 14.5 -3.2 11.2 -3.9 8.9 13.9 0.205 0.043 3.4 1.4
AqE 72 24.0 297.5 -8.4 32.0 -16.7 135 8.2 -26.1 -12.7 -12.0 -7.8 13.5 0.239 0.039 4.3 2.2
AriS 57 20.6 43.5 17.5 159.8 -36.0 142 12.0 -7.1 9.3 0.4 11.0 19.5 0.734 0.03 17.2 13.0
CasE 43 20.7 32.8 65.5 131.0 4.1 288 10.3 -8.3 16.8 -6.5 -4.3 19.8 0.647 0.03 15.3 9.6
CMaN 178 64.9 102.1 -15.1 226.1 -7.5 183 -0.1 -5.3 4.0 -2.1 24.7 9.8 0.214 0.075 2.0 0.5
LeoE 100 51.8 173.5 17.2 238.0 69.5 308 -9.2 -1.5 -5.3 -12.5 13.6 7.2 0.322 0.079 2.9 1.3
LeoC 43 16.2 157.8 14.5 227.0 55.1 271 -9.4 -1.9 -3.3 -11.8 16.8 13.9 0.271 0.032 6.0 2.4
Theia72 254 95.2 102.2 -10.2 221.7 -5.2 278 -1.1 -4.5 4.6 -3.9 28.7 12.3 0.328 0.082 2.8 0.9
Theia72 0 64 24.5 104.7 -13.0 225.3 -4.3 265 -1.7 -4.1 3.6 -4.2 28.3 5.8 0.244 0.06 2.9 1.2
Theia72 1 94 31.4 101.3 -9.3 220.5 -5.6 299 -0.9 -4.3 4.8 -3.8 29.2 8.2 0.294 0.057 3.6 1.7
Theia72 2 95 39.4 101.4 -9.1 220.4 -5.4 266 -0.8 -4.9 5.0 -3.7 28.2 4.9 0.347 0.083 3.0 1.2
OphSE 113 43.6 257.7 -18.4 4.5 12.4 214 -5.1 -11.4 -12.4 -2.4 -14.8 5.3 0.311 0.084 2.6 1.2
ScuN 51 19.8 278.6 -8.1 23.8 0.1 203 0.4 -18.8 -15.9 -8.6 -7.5 15.7 0.851 0.033 18.3 7.0
TOR1A 165 65.6 66.5 13.2 182.1 -24.2 310 4.8 -5.8 11.0 0.3 11.0 20.7 0.476 0.052 6.5 1.9
Theia78 108 40.6 80.2 -4.6 206.5 -22.3 320 0.2 5.2 -6.8 3.9 26.1 11.4 0.31 0.055 4.0 1.8
TOR1B 27 8.9 66.4 15.4 180.2 -22.9 289 3.4 -4.0 7.1 0.2 19.8 10.4 0.514 0.027 13.4 5.2
VulE 85 32.2 309.2 21.6 64.7 -11.5 297 10.0 -5.5 1.8 -15.9 -1.7 7.5 0.314 0.061 3.7 1.4
SCYA-54 301 116.5 16.5 51.3 125.3 -11.4 319 15.5 -6.9 24.0 -9.0 -8.6 14.3 0.418 0.084 3.5 1.5
SCYA-54 0 176 69.1 14.3 51.5 123.9 -11.4 324 15.2 -6.3 23.5 -9.2 -9.0 10.1 0.49 0.077 4.5 1.7
SCYA-54 1 126 47.4 19.6 51.2 127.3 -11.4 311 16.0 -7.8 24.6 -8.8 -8.4 5.8 0.194 0.084 1.6 0.9
SCYA-79 310.8 45.2 84.4 1.7 437 3.7 0.2 5.0 -5.7 -5.8 22.0 0.589
SCYA-79 0 308.9 43.5 82.2 1.8 447 3.2 -0.2 3.7 -5.7 -4.8 11.5 0.536
SCYA-79 1 312.0 46.2 85.7 1.7 431 4.0 0.5 5.8 -5.7 -6.1 12.3 0.313

### v3

@@ 4.1.3 Lithium
The depletion of the Li-6707Å line can be used not just for youth indication as discussed in Section 3.1.5, but also as a powerful tool for age measurement. The EAGLES package provides empirical models of the lithium depletion curve as a function of age, as well functions for computing an age probability posterior for a given stellar Li EW and effective temperature $T_{eff}$. These functions use both the original analytical model (v1; EAGLES23) that we used in Section 3.1.5, and a new machine learning (ML) neural network version (v2; Weaver24).
@@
We multiply the age likelihoods of all remaining stars to produce aggregate posteriors for the entire population, and we do this separately for both the original and ML EAGLES modules. We set the age to the 50th percentile of the posterior, and set the lower and upper uncertainty intervals to the 16th and 84th percentiles, respectively. We exclude ages under 5 Myr, where EAGLES lacks age discrimination and therefore sets the likelihood to its value at 5 Myr (EAGLES23). We only report an upper age limit for populations where the posterior at 5 Myr exceeds half of its peak, which we set to the 95th percentile of the probability posterior. We plot the Lithium depletion age solutions in Figure 8, showing the Li EWs used in the calculation, the best-fit EAGLES Li curves, and the resulting combined age posterior.
The two different EAGLES modules produce broadly similar results. The main difference is in the “Li dip”, which reaches zero about 10 Myr earlier and transitions more sharply analytical version. In populations like AqE, where this dip is well-covered, the analytical model is in close agreement with both the nearly Li-free stars in the dip and the Li-rich stars to the right of it, while the ML version does not capture this sharp transition as closely. However, the ML version seems to produce more consistent results in populations with less coverage, especially populations with a significant Li scatter. The large scatter and Li dip shape in the ML module are produced by an observed scatter in real Li observations (Weaver24). Our harsh quality cuts on binarity and membership may improve this scatter beyond the expectations of the ML model in populations like AqE, but this does not appear to be universal. We therefore include both the analytical and ML results in Figure 8 and Section 4.1.4, but use only the analytical results for membership assessment and quality checks for consistency with previous work (e.g., Kerr24).
@@ 4.1.4 Final Ages
We have produced isochronal ages using several different models, as well as dynamical and lithium depletion ages. The isochronal ages provide a consistent relative age scale derived from the height of the PMS that is available for all populations, unlike the dynamical and lithium depletion methods. However, isochronal ages are often inconsistent with each other and with other age calculation methods in absolute terms (e.g., Herczeg15; Rottensteiner24). The accuracy of these isochronal solutions can be assessed by comparison with the non-isochronal methods.
@@
While agreement with other methods may marginally favor the PARSEC over DSEP-Magnetic, the use of PARSEC is perhaps better motivated by the consistent agreement between that model and our pre-main sequences. In Figure 6, more luminous stars tend to fall below the best-fit DSEP-Magnetic isochrone by up to 0.5 magnitudes, while PARSEC can fit both the upper and lower PMS simultaneously. Among those higher-mass stars, the best-fit ages produced by the DSEP-magnetic models can be up to double the best-fit age for the association as a whole, while PARSEC shows no such discrepancy. The DSEP-Magnetic ages may therefore be biased old in populations with more high-mass PMS stars, making the age results produced by those models a function of not just the height of the PMS, but also the mass function. We therefore conclude that PARSEC produces the most reliable age scale for our analysis.
@@
The ML module, which is designed to better capture natural scatter in these measurements, produces results that align more closely with PARSEC in the populations with high analytical Li ages relative to PARSEC. The apparent high bias in the Li age fits may therefore be driven by coverage that is too incomplete to reject unrealistically old age solutions. Isochrones that do not account for magnetically-driven radius inflation have been shown to produce age results far younger than suggested by Lithium depletion (Jeffries17; Franciosini22), and the BHAC15 isochrone fits appear to exemplify that effect. The PARSEC and DSEP-Magnetic fits, however, both include corrections that modify the mass-radius relationship (Chen14; Feiden16), and our results using those models show no conclusive evidence of systematic age underestimation relative to the Lithium Depletion age.
All dynamical ages are consistent with our PARSEC ages within uncertainties, except in SCYA-79, where the PARSEC ages are poorly constrained. There is no clear evidence of an offset between the dynamical and isochronal ages like what has previously been used to indicate a gas dispersal timescale (e.g., Kerr22a; MiretRoig24). We discuss this point further in Section 5.
@@ 4.2 Demographics
All of the populations we explore in this paper have yet to have their sizes, masses, and total membership comprehensively investigated. Kerr23’s candidate member lists provided the first look at sizes of many of these associations, while some larger populations have also been covered by the (Kounkel19) membership lists. However, these publications did not address all potential sources of bias. Most of these populations represent an understudied demographic of low-mass, isolated populations that has yet to be explored, so detailed demographic work is essential to identify differences and similarities between these populations and the much larger populations that dominate the current literature on young associations.
@@ 4.2.1 Stellar Masses
We compute the masses of individual stars by comparing the stellar photometry to the PARSEC v1.2S isomass track grid from Kerr24, which uses mass sampling every 0.005 $M_{\odot}$ for $0.09M_{\odot}<M<1M_{\odot}$, every 0.01 $M_{\odot}$ between $1.0M_{\odot}<M<2.0M_{\odot}$, every 0.02 $M_{\odot}$ for $2M_{\odot}<M<4M_{\odot}$, and every 0.05 $M_{\odot}$ for $4M_{\odot}<M<20M_{\odot}$. We set the mass to that of the nearest isomass track. Most populations have a maximum stellar mass between 2 and 4 $M_{\odot}$, with no candidates more massive than 4 $M_{\odot}$. These populations therefore lack stars that could constrain the main sequence turn-off ages, which is not surprising given their small total masses (e.g., Chabrier05).
@@ 4.2.2 Correction for Low-Mass Stars
We show a histogram of stars that pass the astrometric and photometric quality flags in Fig. 10, with bins evenly log-sampled between 0.09 and 15 M, and with each object weighed by $P_{fin}$. We smooth the result with a savitsky-golay filter to produce a mass frequency curve, which we also show in Fig. 10, alongside the Chabrier05 IMF scaled to best fit the mass frequency curve for $M>0.25M_{\odot}$ via least-squares optimization. Our mass distributions generally follow the Chabrier05 IMF, although there is some variation, most notably a stellar deficit at $M\sim 0.5$ and an excess at $M\sim 0.25$. This pattern has been shown in several prior publications, and is usually attributed to model innacuracies that displace some stars to lower masses (e.g., Kraus14; Kerr24).
@@
Figure 10: IMF-based mass corrections to all 15 associations where demographics are possible. The dark blue histogram shows the distribution of association members, and a smoothed version of that is shown in light blue. The orange curve shows the Chabrier05 IMF. The missing mass, shown in red, is computed using the volume between the Chabrier05 IMF and the smoothed version of the stellar mass distribution curve. This is used to compute the corrective factors for mass and number ($C_{M,lm}$ and $C_{N,lm}$, respectively), which are annotated on each panel.
All mass frequency curves show fewer stars in the low-mass regime compared to the Chabrier05 IMF, indicating a drop in Gaia completeness there. We take the difference between the mass distribution curve and the scaled IMF to the left of the leftmost crossover point in Figure 10 as the volume missed by Gaia. The corrective factor to number of members is the sum across the entire stellar histogram, plus the sum over the area to the left of the crossover point, divided by the sum over the entire stellar histogram. We limit this calculation to stars with masses $M>0.09M_{\odot}$, or above the PARSEC lower mass limit, which is used as the mass cutoff for inclusion in our stellar sample. The mass corrective factor is computed the same way, but with each histogram and scaled IMF multiplied by the mass at each bin element. We did not bound this calculation in mass, as low-mass and sub-stellar objects can still contribute to the virial state of a stellar population. The resulting corrective factors range from 1.06 to 1.18 in mass and from 1.14 to 1.53 in number. The size of these corrections correlate with distance, with four of five populations with a mass correction less than 1.1 having d<220d<220pc. SCYA-54, which contains the most distant subgroup we cover here, has the largest correction in both mass and number.
@@ 4.2.3 High-mass stars
Some of our mass histograms in Fig. 10 show a deficit of stars with $M\gtrsim 2M_{\odot}$, which may be caused by a lack of quality Gaia astrometry and photometry for stars that saturate ($G\lesssim 3$) (GaiaDR1). However, Hipparcos provides complementary coverage that is essentially complete for stars with $G\lesssim 7$ (Hipparcos97; Hipparcos07). We therefore search each population for Hipparcos stars missed by Gaia. We mark any Hipparcos star as a potential member if its distance to the tenth nearest association member in Kerr23’s space-transverse velocity coordinate space ($d_{10}$) is less than that of the member with the greatest $d_{10}$, following previous SPYGLASS publications. We only consider stars without Gaia astrometry to exclude stars that have already had their membership assessed.
Across all 16 populations, we identify only 5 credible additional members in the Hipparcos catalog across 4 of our associations. Only 3 stars: Mirzam ($\beta$ CMa) and HIP 32848 in CMaN and HIP 4285 in SCYA-54 have masses in the $M\gtrsim 2M_{\odot}$ range where we see a stellar deficit in Fig. 10. We therefore conclude that there is no evidence of our Gaia-based sample systematically excluding high-mass stars, and choose not to apply a broad mass correction. It is possible that high-mass stars are intrinsically more difficult to detect, perhaps due to an increased velocity dispersion produced by the higher binarity rates of high-mass stars (Sullivan21). Follow-up work will therefore be necessary to determine whether some populations truly have mass distributions that differ from a standard IMF.
@@ 4.2.4 Binaries
We estimate the mass contribution from unresolved companions by applying literature binary statistics to our detected primaries. We first identify resolved binaries to avoid double-counting companions when applying binary statistics. Following Kerr24, we search a radius of $10^{4}$ AU on-sky around each star in Gaia. We mark stars with $\Delta v_{T}<3$ km s-1 and $\frac{\Delta\pi}{\pi}<0.2$ as in the same stellar system, with the brightest member marked as the primary, and dimmer members considered companions. These conditions roughly encapsulate the full range of expected velocities for stars near Gaia’s roughly 1 arcsecond separation limit, given typical uncertainties (rizzuto_zodiacal_2018). Unresolved companions move the photometry of a star largely perpendicular to the pre-main sequence, changing its mass estimate relatively little. For measuring demographics, probable unresolved binaries with RUWE>1.2>1.2 can therefore be treated as if they are system primaries without substantially increasing the systematic uncertainties in our association mass estimates.
We follow the methods applied in Kerr25 to estimate the mass in binaries, sampling $10^{4}$ stars from the Chabrier05 IMF at random for $0.09<M<4~M_{\odot}$, and computing the expected mass and number of companions for each sampled star according to the mean mass ratios from Sullivan21 and the companion fractions from Sullivan21 and Duchene13. The expected binary contribution computed from this sampling increases population mass by a factor of 1.23, and increases number of stars by a factor of 1.39. These values produce an average correction to the contribution from binaries for any given stellar primary, including low-mass stars that we correct for in Section 4.2.2.
@@
ASSOCaaUnique choices and properties for specific associations: in CasE, we set stellar membership probabilities below a cut in Fig. 6 to zero to exclude an elevated field sequence. TOR1A and TOR1B are components of the TOR1 association that have velocities inconsistent with common formation. We provide basic position and velocity data for SCYA-79, but do not report population size, mass, or any values deriving from them due to weak separation with the field. SG NbbFor our mass and stellar population measurements, we adopt a 10% systematic uncertainty. MbbFor our mass and stellar population measurements, we adopt a 10% systematic uncertainty. RA Dec l b d $\overline{\mu_{RA}}$ $\overline{\mu_{Dec}}$ $\overline{v_{T,l}}$ $\overline{v_{T,b}}$ $\overline{v_{r}}$ $R_{hm}$ $\sigma_{1D}$ $\sigma_{vir}$ Vir. Rat.
@@
AndS 73 37.0 6.4 30.4 116.3 -32.0 169 14.5 -3.1 11.2 -3.9 8.7 11.1 0.222 0.053 2.9 1.3
AqE 80 33.1 297.5 -8.7 31.7 -16.8 135 8.3 -26.0 -12.6 -12.0 -7.8 12.2 0.245 0.048 3.6 1.8
AriS 70 30.7 43.5 17.6 159.8 -35.9 142 12.1 -7.2 9.4 0.5 10.7 18.0 0.782 0.038 14.4 10.5
CasE 53 27.5 32.6 65.6 130.9 4.1 289 10.3 -8.2 16.9 -6.4 -4.2 21.8 0.704 0.033 15.1 9.9
CMaN 191 73.0 102.1 -15.1 226.1 -7.5 183 -0.1 -5.3 4.0 -2.1 24.6 8.6 0.209 0.086 1.7 0.5
LeoE 100 51.8 173.5 17.2 238.0 69.5 308 -9.2 -1.5 -5.3 -12.5 13.6 6.9 0.323 0.08 2.8 1.2
LeoC 43 16.2 157.8 14.5 227.0 55.1 271 -9.4 -1.9 -3.3 -11.8 16.8 14.2 0.246 0.031 5.5 1.9
Theia72 273 107.8 102.2 -10.2 221.7 -5.2 278 -1.1 -4.5 4.6 -3.9 28.6 12.0 0.33 0.088 2.7 0.8
Theia72 0 68 26.4 104.6 -12.9 225.3 -4.4 265 -1.7 -4.1 3.7 -4.2 28.3 5.7 0.255 0.063 2.9 1.1
Theia72 1 100 34.4 101.4 -9.3 220.5 -5.5 298 -0.9 -4.3 4.9 -3.8 29.2 7.7 0.289 0.062 3.3 1.5
Theia72 2 105 47.0 101.4 -9.2 220.4 -5.5 266 -0.8 -4.9 5.0 -3.7 28.2 5.3 0.312 0.088 2.5 1.0
OphSE 127 55.0 257.7 -18.4 4.5 12.4 213 -5.1 -11.4 -12.4 -2.3 -14.7 4.9 0.326 0.098 2.3 1.0
ScuN 60 22.8 278.3 -8.2 23.6 0.3 205 0.3 -18.7 -15.9 -8.6 -7.5 14.0 0.81 0.037 15.3 6.1
TOR1A 185 79.1 66.4 13.3 182.0 -24.2 311 4.8 -5.7 11.0 0.3 10.8 17.5 0.475 0.062 5.4 1.7
Theia78 120 49.9 80.3 -4.6 206.6 -22.2 320 0.2 5.2 -6.8 3.9 26.3 11.2 0.309 0.062 3.5 1.6
TOR1B 34 16.2 66.2 15.5 180.0 -22.9 290 3.3 -3.9 7.1 0.2 21.0 8.2 0.43 0.041 7.3 3.0
VulE 89 36.1 309.2 21.6 64.7 -11.5 297 10.0 -5.5 1.8 -15.9 -1.7 8.1 0.301 0.062 3.5 1.3
SCYA-54 306 119.4 16.5 51.4 125.3 -11.4 319 15.5 -6.9 24.0 -9.0 -8.8 14.4 0.414 0.084 3.5 1.5
SCYA-54 0 177 70.1 14.3 51.5 123.9 -11.3 324 15.2 -6.3 23.6 -9.2 -9.0 10.3 0.485 0.076 4.5 1.6
SCYA-54 1 129 49.3 19.6 51.2 127.3 -11.4 311 16.0 -7.8 24.6 -8.8 -8.5 6.0 0.216 0.084 1.8 1.0
SCYA-79 310.7 45.1 84.3 1.7 438 3.7 0.2 5.0 -5.8 -5.9 22.5 0.581
SCYA-79 0 308.8 43.5 82.2 1.8 447 3.2 -0.2 3.7 -5.8 -4.9 11.6 0.504
SCYA-79 1 312.0 46.2 85.7 1.7 431 4.0 0.5 5.8 -5.8 -6.1 12.7 0.309